OpenStudy (dimples97):

could someone please show me how to simplify this expression? will give medal thanks! 3x(4x^4 – 5x)

OpenStudy (jtug6):

Okay so you can already see theres a 3x sticking on the outside of the parenthesis, so to simplify we should do what to the 3x and the two terms (4x^4 and - 5x)

OpenStudy (legomyego180):

\[3x(4x^4-5x)\] So you could rewrite this by distributing the 3x, or by factoring an x out of the parentheeses. Im not sure which your instructor wants.

OpenStudy (jtug6):

Oh I misread actually they want you to simplify not expand xD. Ok so scratch what I just said (sorry I'm sleepy). Look for the common variable in the parenthesis (x) and factor it out. ^^

OpenStudy (jtug6):

Right I think it's the factoring part since they said simplify thats why I was confused haha

OpenStudy (legomyego180):

So when you factor you just take an x term out of the parentheeses. \[(4x^4-5x) = x(4x^3-5)\] Because the x is outside of the parentheeses we can multiply it by the 3x now: \[(3x)(x)(4x^3-5)=3x^2(4x^3-5)\]

OpenStudy (legomyego180):

You could also expand this out and simplify it further by writing it as: \[12x^5-15x^2\]